The cost of retained earnings equity for KRAS company is 7.5 percent, when the expected (risk free) return on U.S. treasury bills is 3 percent, the firm's beta is 1.5 and the market return is ??
Answer #1

Cost of Equity {Ke} = 7.5% Risk - free rate { Rf} = 3.1. Beta = 1.5 - Ke = Rp + Beta x [Rm - Rf] 7.5 = 3 + 1.5 x [ RM - 3] 7.